A 0.240g of a compound of oxygen and element X was found by analysis to contain 0.960g of X and 0.144g of oxygen

1. Calculate the percentage composition by weight

2. Calculate from the above data, three(3) possible weight for X relative to oxygen








1
Expert's answer
2022-05-20T02:12:02-0400

comment: The mass of x is slightly incorrect in the case condition (0,96). I did the problem by counting it as 0,096.


1) To find the percentage by mass, we divide the mass of oxygen or x by the total mass and muuultiply by 100.

for oxygen : 0.144 / 0.24 * 100 = 60%

for X : 0.096 / 0.24 * 100 = 40%

2) 0.144 g of oxygen corresponds to 0.096 g of X. so, x is 1.5 times less than oxygen. that is, 0.144 / 0.096 = 1.5

Find how much x corresponds to 3 oxygen atoms (16 * 3 = 48 g). To do this, we divide the mass of 3 oxygen atoms by 1.5:

48 / 1.5 = 32gr (this may be sulfur)

The approximate formula of the substance is SO3

ANSWER:

1. 60% oxygen 40% X

2. 32gr

The approximate formula of the substance is SO3